# Doped resonating valence bond states: How robust are the spin ice phases in 3D Rydberg arrays

### v1

@@ III Cubic lattice
Then we consider another three-dimensional general RK model where each atom is positioned on the link of cubic lattice. This setting is the generalization of the Rydberg implementation of the two-dimensional quantum square ice model into the three-dimensional cubic lattice [63]. It’s worth noting that, unlike the diamond lattice general RK model, this model is not easy to realize through Rydberg arrays, but its physics is similar as the diamond one. Due to the simpler lattice structure, more physics can be explored here. Firstlt, we go back to the Hamiltonian of pure RK model in this setup, which is defined as follows:
@@
where the summation is over all plaquettes within the cubic lattice. The first term is the kinetic term which represents ring-exchange process on flippable plaquettes, while the second term counts the total number of flippable plaquettes. At the RK point $V=t$, the ground state is exactly solvable as described by equal-weight superposition of all dimer configurations within each topological sector.
We evaluate the dispersion of cubic lattice QDM at the RK point $V=t=1$ based on SMA, with the computational approach similar to that described in Sec. II. The system size of the cubic lattice in our simulation is $32\times 32\times 32$. Considering the cubic symmetry of the lattice structure, without loss of generality we only consider the correlations of dimer density along the $\overrightarrow{x}$ direction $D_{x}$.
@@
Figure 8: Dispersion of dimer of the cubic lattice at the RK point $V=t=1$. Under the local constraint of one-dimer-per-site, the dispersion of the QDM on a cubic lattice based on the SMA along the path $\Gamma\rightarrow(\pi,0,0)\rightarrow(\pi,\pi,0)\rightarrow(\pi,\pi,\pi)\rightarrow\Gamma$.
@@
where the summation of third and the fourth terms are performed throughout all links, the third term is the dimer hopping term that incorporates hopping of monomers between different sites, and the fourth term counts the total number of hoppable dimers.
@@
This symmetry divides the overall Hilbert space into different sectors characterized by their total dimer occupation $M$. Within each sector, the ground state is also exactly soluble at the generalized RK point $t=V$ and $t^{\prime}=V^{\prime}$, with energies being zero, and the wave functions described by equal-weight superposition of all configurations within the sector. Without loss of generality, in the following we only present results at a specific point $t=V=t^{\prime}=V^{\prime}=1$. As discussed previously, monomers are defined respect to the closest background configurations that satisfy local constraints. For example, if the doping density of one-dimer-per-site configurations exceeds 50%, we adjust the background to configurations with two dimers per site to define the monomers.
@@
To investigate the impact of monomer doping in the quantum dynamical behaviors, we calculate the dispersions of dynamical dimer density correlations within two scenarios: the strictly one-dimer-per-site manifold (without monomers) and manifolds doped with monomers at densities of 1.22%, 25% and 49.8%. These calculations involve sampling connected configurations and applying statistical averaging to approximate the system’s dispersion, and the result is shown in Fig. 9. Without monomer doping, there exhibits a gapless mode with quadratic dispersion at $(\pi,\pi,\pi)$ that corresponds to the “photon” excitations of the $U(1)$ Coulomb liquid. This photon mode is gapped out upon monomer doping, coincides with that observed in the pyrochlore lattice, and can be interpreted as the condensation of monomers that Higgses the $U(1)$ gauge field, destroys the deconfined Coulomb phase and results a trivial confined phase.
@@
Figure 9: Dynamical dimer density dispersion on a cubic lattice with doping along the Brillouin zone path $\Gamma\rightarrow(\pi,0,0)\rightarrow(\pi,\pi,0)\rightarrow(\pi,\pi,\pi)\rightarrow\Gamma$. The different colored circles correspond to the doping densities $1.22\%,25\%$ and $49.8\%$ in one-dimer-per-site.
@@
Figure 10: Strings in cubic lattice. The model illustrates one of the configurations of a two-dimer per site within a dimer model of size $L=2$ that satisfies periodic boundary conditions. In the three-dimensional dimer model, the string is a cube with edge length $(l=1)$, surrounding the lattice site. Some dimers are intersected perpendicularly by the faces of the string. The direct lattice vectors in the $\overrightarrow{x},\overrightarrow{y}$ and $\overrightarrow{z}$ directions are set to have a unit length of 1.
@@
The fate of the Coulomb liquids upon doping can be also examined by probing the non-local “surface” operators, which serve as a diagnostic for quantum spin liquids (QSLs). Here, we focus on the diagonal operator $Z=\prod_{i\in s}\sigma_{i}^{z}$, where $\prod_{i\in s}$ denotes the product of dimers that intersect with the surface $s$, $\sigma_{i}^{z}=1-2n_{i}$, $n_{i}=0$ or $1$ respectively indicates whether there is a dimer at position $i$, see Fig. 10. This non-local operator $Z$, that measures the parity of the dimers intersecting with the surface, characterizes how well the local dimer constraint is satisfied. For example, for the smallest cubic surface that encloses only a single site of the cubic lattice (Fig.  10), $\langle Z\rangle=\pm 1$ holds for any undoped dimer configurations, where the sign depends on the parity of the number of dimers attached to each site. For generic surfaces, $\langle Z\rangle=1$ holds for configurations that each site touches even number of dimers, while $\langle Z\rangle=(-1)^{\textrm{\# of enclosed sites}}$ for configurations that each site touches odd number of dimers. In presence of monomer doping, the local constraints are no longer strictly satisfied, leading to a decrease in $|\langle Z\rangle|$. Also note that in two-dimensional lattices, such surfaces corresponds to closed strings, hence the definition of $Z$ devolves into that in Ref. [9].
[…]
To figure out whether infinitesimal amount of monomer defects could destroy the structure of the $U(1)$ Coulomb liquid, we choose a configuration with $0.024414\%$ monomer doping on the background of two dimer per site, and examine the evolution of string operator expectation value $\langle Z\rangle$ with increasing surface size $l$. As shown in Fig. 11(b), we find a superexponential decay of $\langle Z\rangle$ with increasing $l$, suggesting that the $U(1)$ Coulomb liquid is unstable against infinitesimal monomers in the thermodynamic limit. Nevertheless, in cold-atom experiments where the atom arrays are prepared with limited system sizes, we still observe a pronounced $\langle Z\rangle$ when monomer doping is small. Hence we still expect to observe Coulomb liquid physics in cold-atom platform even if small amount of monomer defects are presented.
@@
Figure 11: (a) String operator as a function of the doping density. (b) String operator as a function of the liner sizes of string for monomer density 0.024414% (black dashed line in (a)).
@@
Figure 12: Dimer off-diagonal correlation at different doping density. The circles represent the off-diagonal correlations along the $\overrightarrow{x}$ for the different distances $|i-j|$ in a cubic lattice. Different colors indicate varying doping densities ($0\%,1.5625\%,17.1875\%,34.375\%,48.4375\%$ and $84.375\%$) in (a) one-dimer-per-site and (b) two-dimers-per-site configurations. The distance between two neighboring links is set to 1
@@
We have also evaluated the dimer off-diagonal correlations to probe the possibility of proximate long-range order. The detailed calculation scheme is similar to that described in Sec. II, and the results are shown in Fig. 12. In absence of monomer defects, we find vanishing off-diagonal dimer correlations as is forbidden by the local constraints, similar to the case of the diamond lattice. As the local constraint become weakened through the introduction of monomers, the system immediately establishes a finite off-diagonal long-range correlations as a consequence of monomer condensation.
@@
Figure 13: Dispersion of monomer density correlation along the paths $\Gamma\rightarrow(\pi,0,0)\rightarrow(\pi,\pi,0)\rightarrow(\pi,\pi,\pi)\rightarrow\Gamma$ for configurations based on the one-dimer-per-site (a) and two-dimers-per-site (b), with doping density of $1.221\%$, $25\%$, 49.9878% and 70.0073%.
@@
We also evaluated the dispersion of monomer density correlations in the cubic lattice along the path $\Gamma\rightarrow\left(\pi,0,0\right)\rightarrow\left(\pi,\pi,0\right)\rightarrow\left(\pi,\pi,\pi\right)\rightarrow\Gamma$ based on method described in II. In Fig. 13, we present the results based on one-dimer-per-site and two-dimers-per-site configurations with doping density ($1.221\%$, $25\%$, $49.9878\%$ and $70.0073\%$). In our SMA calculations, we observe that the structure factor in the denominator vanishes at the $\left(\pi,0,0\right)$, $\left(\pi,\pi,0\right)$ and $\left(\pi,\pi,\pi\right)$ points, making the results ambiguous. To further investigate the nature of excitations at these points, we perform a finite-size scaling analysis, examining the asymptotic behavior of the dispersions at the nearby points $\left(\frac{2\pi}{L}\left(\frac{L}{2}-1\right),0,0\right)$, $\left(\pi,\frac{2\pi}{L}\left(\frac{L}{2}-1\right),0\right)$ and $(\pi,\pi,\frac{2\pi}{L}(\frac{L}{2}-1))$ at different system sizes $L\times L\times L$. As shown in Fig. 14, we find that the gap extrapolates to zero at the $\left(\pi,0,0\right)$ for both doped one-dimer-per-site and two-dimers-per-site states, and $\left(\pi,\pi,0\right)$ for only two-dimers-per-site state with increasing system size, almost independent of the monomer density. Fig. 15 displays that the gap at $\left(\pi,\pi,0\right)$ for one-dimer-per-site state will open after doping. In contrast, the dispersion at $\left(\pi,\pi,\pi\right)$ extrapolates to some finite values and increases upon monomer doping. Therefore, we conclude that in presence of monomer doping, excitations at the $\Gamma$, $\left(\pi,0,0\right)$ and $\left(\pi,\pi,0\right)$ points remain gapless while excitations at $\left(\pi,\pi,\pi\right)$ becomes gapped out.
@@
Figure 14: The gap of monomers is shown as a function of the inverse squared system size for (a) one-dimer-per-site and (b) two-dimers-per-site doped with $1\%$ and $25\%$ monomers, respectively, at $\left(\frac{2\pi}{L}\left(\frac{L}{2}-1\right),0,0\right)$, $\left(\pi,\frac{2\pi}{L}\left(\frac{L}{2}-1\right),0\right)$ and $(\pi,\pi,\frac{2\pi}{L}(\frac{L}{2}-1))$.
@@ IV Discussion and outlook
Motivated by the recent developments on Rydberg atom arrays, in this work we investigate the relevant extended QDMs on three-dimensional bipartite lattice geometries with the introduction of monomers. For both the diamond and the cubic lattices, we find that even a small amount of monomer doping disrupts the deconfined $U(1)$ Coulomb liquid and stabilizes off-diagonal long-range order. However, in a finite size (as in cold-atom experiment), the property of QSI will be kept in a certain region like a crossover after doping.
Our results on the extended QDM have significant implications for experimental realizations, particularly in cold atom setups. One of the basic assumption in our model study is that monomer defects are condensed and enters the ground state wave functions upon doping. In practical experiments, monomers are often controlled by some effective chemical potential. If monomers remain gapped, they can be integrated out, leaving the $U(1)$ Coulomb phase unaffected. However, if the monomer gap close and condensation occur, this would lead to the Higgsing of the $U(1)$ gauge field, resulting in a trivial phase with off-diagonal long-range order in the thermodynamic limit. Despite this, real cold atom experiments are typically constrained by finite lattice sizes, which means that the system may not reach the thermodynamic limit. Consequently, we still expect $U(1)$ Coulomb liquid behavior within these experimental conditions if the doping is small.
@@ Acknowledgements.
We thank the helpful discussions with Yang Qi, Fabien Alet, Sylvain Capponi and Pranay Patil. Y.C.W. acknowledges the support from Zhejiang Provincial Natural Science Foundation of China (Grant No. LZ23A040003), and the support from the High-Performance Computing Centre of Hangzhou International Innovation Institute of Beihang University. The work is supported by the start-up funding of Westlake University. The authors thank the high-performance computing center of Westlake University and Beijng PARATERA Tech Co.,Ltd. for providing HPC resources.

### v2

@@ III Cubic lattice
Then we consider another three-dimensional general RK model where each atom is positioned on the link of cubic lattice. This setting is the generalization of the Rydberg implementation of the two-dimensional quantum square ice model into the three-dimensional cubic lattice [65]. It’s worth noting that, unlike the diamond lattice general RK model, this model is not easy to realize through Rydberg arrays, but its physics is similar to the diamond one. Due to the simpler lattice structure, more physics can be explored here. First, we go back to the Hamiltonian of pure RK model in this setup, which is defined as follows:
@@
where the summation is over all plaquettes within the cubic lattice. The first term is the kinetic term which represents the ring-exchange process on flippable plaquettes, while the second term counts the total number of flippable plaquettes. At the RK point $V=t$, the ground state is exactly solvable as described by equal-weight superposition of all dimer configurations within each topological sector.
We evaluate the dispersion of cubic lattice QDM at the RK point $V=t=1$ based on SMA, with the computational approach similar to that described in Sec. II. The system size of the cubic lattice in our simulation is $32\times 32\times 32$. Considering the cubic symmetry of the lattice structure, without loss of generality, we only consider the correlations of dimer density along the $\overrightarrow{x}$ direction $D_{x}$.
@@
Figure 8: Dispersion of dimer of the cubic lattice at the RK point $V=t=1$. Under the local constraint of one dimer per site, the dispersion of the QDM on a cubic lattice based on the SMA along the path $\Gamma\rightarrow(\pi,0,0)\rightarrow(\pi,\pi,0)\rightarrow(\pi,\pi,\pi)\rightarrow\Gamma$.
@@
where the summation of the third and the fourth terms are performed throughout all links, the third term is the dimer hopping term that incorporates hopping of monomers between different sites, and the fourth term counts the total number of hoppable dimers.
@@
This symmetry divides the overall Hilbert space into different sectors characterized by their total dimer occupation $M$. Within each sector, the ground state is also exactly soluble at the generalized RK point $t=V$ and $t^{\prime}=V^{\prime}$, with energies being zero, and the wave functions described by equal-weight superposition of all configurations within the sector. Without loss of generality, in the following we only present results at a specific point $t=V=t^{\prime}=V^{\prime}=1$. As discussed previously, monomers are defined with respect to the closest background configurations that satisfy local constraints. For example, if the doping density of one-dimer-per-site configurations exceeds 50%, we adjust the background to configurations with two dimers per site to define the monomers.
@@
To investigate the impact of monomer doping in the quantum dynamical behaviors, we calculate the dispersions of dynamical dimer density correlations within two scenarios: the strictly one-dimer-per-site manifold (without monomers) and manifolds doped with monomers at densities of 1.22%, 25.0%, and 49.8%. These calculations involve sampling connected configurations and applying statistical averaging to approximate the system’s dispersion, and the result is shown in Fig. 9. Without monomer doping, there exhibits a gapless mode with quadratic dispersion at $(\pi,\pi,\pi)$ that corresponds to the “photon” excitations of the U(1) Coulomb liquid. This photon mode is gapped out upon monomer doping, coincides with that observed in the pyrochlore lattice, and can be interpreted as the condensation of monomers that Higgses the U(1) gauge field, destroys the deconfined Coulomb phase and results a trivial confined phase.
@@
Figure 9: Dynamical dimer density dispersion on a cubic lattice with doping along the Brillouin zone path $\Gamma\rightarrow(\pi,0,0)\rightarrow(\pi,\pi,0)\rightarrow(\pi,\pi,\pi)\rightarrow\Gamma$. The different colored circles correspond to the doping densities $1.22\%,25.0\%$ and $49.8\%$ in one dimer per site.
@@
Figure 10: Surface in cubic lattice. The model illustrates one of the configurations of a two-dimer per site within a dimer model of size $L=2$ that satisfies periodic boundary conditions. In the three-dimensional dimer model, the surface is a cube with edge length $(l=1)$, surrounding the lattice site. Some dimers are intersected perpendicularly by the surface. The direct lattice vectors in the $\overrightarrow{x},\overrightarrow{y}$ and $\overrightarrow{z}$ directions are set to have a unit length of 1.
@@
The fate of the Coulomb liquids upon doping can be also examined by probing the non-local “surface” operators, which serve as a diagnostic for quantum spin liquids (QSLs). Here, we focus on the diagonal operator $Z=\prod_{i\in s}\sigma_{i}^{z}$, where $\prod_{i\in s}$ denotes the product of dimers that intersect with the surface $s$, $\sigma_{i}^{z}=1-2n_{i}$, $n_{i}=0$ or $1$ respectively indicates whether there is a dimer at position $i$, see Fig. 10. This non-local operator $Z$, which measures the parity of the dimers intersecting with the surface, characterizes how well the local dimer constraint is satisfied. For example, for the smallest cubic surface that encloses only a single site of the cubic lattice (Fig. 10), $\langle Z\rangle=\pm 1$ holds for any undoped dimer configurations, where the sign depends on the parity of the number of dimers attached to each site. For generic surfaces, $\langle Z\rangle=1$ holds for configurations that each site touches even number of dimers, while $\langle Z\rangle=(-1)^{\textrm{\# of enclosed sites}}$ for configurations that each site touches odd number of dimers. In presence of monomer doping, the local constraints are no longer strictly satisfied, leading to a decrease in $|\langle Z\rangle|$. Also note that in two-dimensional lattices, such surfaces correspond to closed strings, hence the definition of $Z$ devolves into that in Ref. [10]. Due to the complexity of the diamond lattice structure, the actual calculation of this physical quantity is rather complicated. Considering that the physics is the same as in the cubic lattice, we have therefore calculated the surface operator only for the cubic lattice.
@@
Figure 11: (a) Surface operator as a function of the doping density. (b) Surface operator as a function of the liner sizes for monomer density 0.0244% (black dashed line in (a)).
We consider doping based on background configurations of two dimers per site. In Fig. 11(a), we plot the expectation value of the non-local operator $\langle Z\rangle$ versus the monomer density for different surface sizes with linear dimension $(l=3,5,7,9,11,13$ and $15)$. At the generalized RK point, the expectation value of the non-local surface operator $\langle Z\rangle$ can be calculated as $\langle Z\rangle=\frac{1}{N_{C}}\sum_{C}(-1)^{m(C)}$, where $(-1)^{m(C)}$ measures the parity of the dimers intersecting with the surface for a given configuration $C$, and $N_{C}$ is the total number of configurations $C$ connected by local dimer moves in the doped QDM Eq. (13). As shown in Fig. 11(a), we find that with increasing monomer doping, $\langle Z\rangle$ rapidly decreases to 0. In addition, with the increase as of the surface size $l$, $\langle Z\rangle$ also rapidly decreases.
To figure out whether infinitesimal amount of monomer defects could destroy the structure of the U(1) Coulomb liquid, we choose a configuration with $0.024414\%$ monomer doping on the background of two dimer per site, and examine the evolution of surface operator expectation value $\langle Z\rangle$ with increasing surface size $l$. As shown in Fig. 11(b), we find a superexponential decay of $\langle Z\rangle$ with increasing $l$, suggesting that the U(1) Coulomb liquid is unstable against infinitesimal monomers in the thermodynamic limit. Nevertheless, in cold-atom experiments where the atom arrays are prepared with limited system sizes, we still observe a pronounced $\langle Z\rangle$ when monomer doping is small. Hence we still expect to observe Coulomb liquid physics in cold-atom platform even if a small amount of monomer defects are presented.
Figure 12: Stability phase diagram of the dimer model on the cubic lattice. The horizontal axis represents the doping density and the vertical axis denotes the system size ($L\times L\times L$). The color map indicates the expectation value of the surface operator. The surface size is L/21L/2-1
@@
Figure 13: Dimer off-diagonal correlation at different doping density. The circles represent the off-diagonal correlations along the $\overrightarrow{x}$ for the different distances |ij||i-j| in a cubic lattice. Different colors indicate varying doping densities ($0\%,1.56\%,17.2\%,34.4\%,48.4\%$ and $84.4\%$) in (a) one-dimer-per-site and (b) two-dimers-per-site configurations. The distance between two neighboring links is set to 1
To clearly illustrate the crossover between effective stable states in small system sizes, we present in Fig. 12 the “stability phase diagram” of the model. We calculate the expectation value of a three-dimensional surface operator $\langle Z\rangle$ of scale l=L/21l=L/2-1 for systems of different sizes ($L\times L\times L$, $L=8,16,24,32$). Under the background of two dimers per site, the color map of $\langle Z\rangle$ is used to characterize the effective U(1) Coulomb liquid at different doping densities. As shown in Fig. 12, the results are fully consistent with the preceding analysis: the characteristic signal of the U(1) Coulomb liquid is extremely sensitive to the size of the system. However, in finite systems this signal still emerges within a very narrow region of doping density.
@@
We have also evaluated the dimer off-diagonal correlations to probe the possibility of proximate long-range order. The detailed calculation scheme is similar to that described in Sec. II, and the results are shown in Fig. 13. In the absence of monomer defects, we find vanishing off-diagonal dimer correlations as is forbidden by the local constraints, similar to the case of the diamond lattice. As the local constraint becomes weakened through the introduction of monomers, the system immediately establishes a finite off-diagonal long-range correlation as a consequence of monomer condensation.
@@
Figure 14: Dispersion of monomer density correlation along the paths $\Gamma\rightarrow(\pi,0,0)\rightarrow(\pi,\pi,0)\rightarrow(\pi,\pi,\pi)\rightarrow\Gamma$ for configurations based on the one dimer per site (a) and two dimers per site (b), with doping density of $1.22\%$, $25.0\%$, 49.9% and 70.0%.
@@
We also evaluated the dispersion of monomer density correlations in the cubic lattice along the path $\Gamma\rightarrow\left(\pi,0,0\right)\rightarrow\left(\pi,\pi,0\right)\rightarrow\left(\pi,\pi,\pi\right)\rightarrow\Gamma$ based on method described in II. In Fig. 14, we present the results based on one-dimer-per-site and two-dimers-per-site configurations with doping density ($1.22\%$, $25.0\%$, $49.9\%$ and $70.0\%$). In our SMA calculations, we observed that at the momentum points $\left(\pi,0,0\right)$, $\left(\pi,\pi,0\right)$ and $\left(\pi,\pi,\pi\right)$, both the numerator and the denominator (the structure factor) of $\omega_{SMA}$ approach zero, resulting in an indeterminate ’$0/0$’ form. This makes it impossible to directly determine whether $\omega_{SMA}$ vanishes at these points. To further investigate the nature of excitations at these points, we perform a finite-size scaling analysis, examining the asymptotic behavior of the dispersions at the nearby points $\left(\frac{2\pi}{L}\left(\frac{L}{2}-1\right),0,0\right)$, $\left(\pi,\frac{2\pi}{L}\left(\frac{L}{2}-1\right),0\right)$ and $(\pi,\pi,\frac{2\pi}{L}(\frac{L}{2}-1))$ at different system sizes $L\times L\times L$. As $L\to\infty$, these adjacent momenta point approach the singular points. As shown in Fig. 15, we find that the gap extrapolates to zero at the $\left(\pi,0,0\right)$ for both doped one-dimer-per-site and two-dimers-per-site states, and $\left(\pi,\pi,0\right)$ for only two-dimers-per-site state with increasing system size, almost independent of the monomer density. Fig. 16 displays that the gap at $\left(\pi,\pi,0\right)$ for one-dimer-per-site state will open after doping. In contrast, the dispersion at $\left(\pi,\pi,\pi\right)$ extrapolates to some finite values and increases upon monomer doping. Therefore, we conclude that in presence of monomer doping, excitations at the $\Gamma$, $\left(\pi,0,0\right)$ and $\left(\pi,\pi,0\right)$ points remain gapless while excitations at $\left(\pi,\pi,\pi\right)$ becomes gapped out.
@@
Figure 15: Finite size scaling of gap at ($\pi$,0,0), ($\pi$$\pi$,0) and ($\pi$$\pi$$\pi$). The gap of monomers is shown as a function of the inverse squared system size for (a) one dimer per site and (b) two dimers per site doped with $1\%$ and $25\%$ monomers, respectively, at $\left(\frac{2\pi}{L}\left(\frac{L}{2}-1\right),0,0\right)$, $\left(\pi,\frac{2\pi}{L}\left(\frac{L}{2}-1\right),0\right)$ and $(\pi,\pi,\frac{2\pi}{L}(\frac{L}{2}-1))$.
@@ IV Discussion and outlook
Motivated by the recent developments on Rydberg atom arrays, in this work we investigate the relevant extended QDMs on three-dimensional bipartite lattice geometries with the introduction of monomers. For both diamond and cubic lattices, we find that even a small amount of monomer doping disrupts the deconfined U(1) QSI and stabilizes off-diagonal long-range order. However, in a finite size (as in cold-atom experiment), the property of QSI will be kept in a certain region like a crossover after doping.
Our results on the extended QDM have significant implications for experimental realizations, particularly in cold atom setups. One basic assumption in our model study is that monomer defects are condensed and enter the ground state wave functions upon doping. In practical experiments, monomers are often controlled by some effective chemical potential. If monomers remain gapped, they can be integrated out, leaving the U(1) Coulomb phase unaffected. However, if the monomer gap close and condensation occur, this would lead to the Higgsing of the U(1) gauge field, resulting in a trivial phase with off-diagonal long-range order in the thermodynamic limit. Despite this, real cold atom experiments are typically constrained by finite lattice sizes, which means that the system may not reach the thermodynamic limit. Consequently, we still expect the behavior of the U(1) Coulomb liquid to be in these experimental conditions if the doping is small.
@@ Acknowledgements.
We thank the helpful discussions with Yang Qi, Fabien Alet, Sylvain Capponi and Pranay Patil. Y.C.W. and C.L acknowledges the support from the Natural Science Foundation of China (Grant No. 12474216 and 12564021) and Zhejiang Provincial Natural Science Foundation of China (Grant No. LZ23A040003), and the support from the High-Performance Computing Centre of Hangzhou International Innovation Institute of Beihang University. The work is supported by the the Scientific Research Project (No.WU2024B027) and start-up funding of Westlake University. The authors thank the high-performance computing center of Westlake University and Beijng PARATERA Tech Co.,Ltd. for providing HPC resources.